Two point charges are placed as follows: charge q1= -1.50 nC is at y= +6.00 m and charge q2= +3.20 nC is at the origin. What is the total force (magnitude and direction) exerted by these two charges on a negative point charge q3= -5.00 nC located at (2.00 m, -4.00 m)?


1
Expert's answer
2021-07-19T14:24:28-0400

Let's first find the magnitude of the force with which the charge "q_1" acts on charge "q_3":


"F_{13}=\\dfrac{kq_1q_3}{r_{13}^2}=\\dfrac{kq_1q_3}{(\\sqrt{r_{12}^2+r_{23}^2})^2}\n\u200b\n \n\u200b\n ,""F_{13}=\dfrac{9\cdot10^9\ \dfrac{Nm^2}{C^2}\cdot1.5\cdot10^{-9}\ C\cdot5\cdot10^{-9}\ C}{(\sqrt{(6\ m)^2+(\sqrt{(2\ m)^2+(4\ m)^2})^2})^2}=1.205\cdot10^{-9}\ N."

Similarly, let's find the magnitude of the force with which the charge "q_2" acts on charge "q_3":


"F_{23}=\dfrac{kq_2q_3}{r_{23}^2},""F_{23}=\dfrac{9\cdot10^9\ \dfrac{Nm^2}{C^2}\cdot3.2\cdot10^{-9}\ C\cdot5\cdot10^{-9}\ C}{(\sqrt{(2\ m)^2+(4\ m)^2})^2}=7.2\cdot10^{-9}\ N."

Let's find the projections of forces "F_{13}" and "F_{23}" on axis "x" and "y":


"F_{13,x}=F_{13}\cdot\dfrac{r_{23}}{r_{13}}=1.205\cdot10^{-9}\ N\cdot\dfrac{\sqrt{(2\ m)^2+(4\ m)^2}}{\sqrt{(6\ m)^2+(\sqrt{(2\ m)^2+(4\ m)^2})^2}},""F_{13,x}=7.2\cdot10^{-10}\ N.""F_{13,y}=-F_{13}\cdot\dfrac{r_{12}}{r_{13}}=-1.205\cdot10^{-9}\ N\cdot\dfrac{6\ m}{\sqrt{(6\ m)^2+(\sqrt{(2\ m)^2+(4\ m)^2})^2}},""F_{13,y}=-9.66\cdot10^{-10}\ N.""F_{23, x}=-F_{23}\cdot\dfrac{r_{23}}{r_{13}}=-7.2\cdot10^{-9}\ N\cdot\dfrac{\sqrt{(2\ m)^2+(4\ m)^2}}{\sqrt{(6\ m)^2+(\sqrt{(2\ m)^2+(4\ m)^2})^2}},""F_{23, x}=-4.3\cdot10^{-9}\ N.""F_{23,y}=F_{23}\cdot\dfrac{r_{12}}{r_{13}}=7.2\cdot10^{-9}\ N\cdot\dfrac{6\ m}{\sqrt{(6\ m)^2+(\sqrt{(2\ m)^2+(4\ m)^2})^2}},""F_{23,y}=5.77\cdot10^{-9}\ N."

Then, we get:


"F_x=F_{13,x}+F_{23,x}=7.2\cdot10^{-10}\ N+(-4.3\cdot10^{-9}\ N)=-3.58\cdot10^{-9}\ N,""F_y=F_{13,y}+F_{23,y}=(-9.66\cdot10^{-10}\ N)+5.77\cdot10^{-9}\ N=4.8\cdot10^{-9}\ N."

We can find the total electrostatic force on the charge "q_3" from the Pythagorean theorem:


"F=\sqrt{F_x^2+F_y^2},""F=\sqrt{(-3.58\cdot10^{-9}\ N)^2+(4.8\cdot10^{-9}\ N)^2}=5.98\cdot10^{-9}\ N."

We can find the direction of the total electrostatic force on the charge "q_3" from the geometry:


"tan\theta=\dfrac{F_y}{F_x},""\theta=tan^{-1}(\dfrac{F_y}{F_x}),""\theta=tan^{-1}(\dfrac{4.8\cdot10^{-9}\ N}{-3.58\cdot10^{-9}\ N})=53.3^{\circ}."

The total electrostatic force on the charge "q_3" directed at "53.3^{\circ}" below the "x"-axis.